OpenStudy (jmansedg):

Using the Pythagorean theorem, find the measure of the hypotenuse when a = 4ft and b = 6ft. A. c = 7ft B. c = 7.2ft C. c = 10ft D. c = 10.2ft

OpenStudy (anonymous):

it would be B

OpenStudy (thecalchater):

a^2+b^2=c^2 so plug in a and b and what do you get?

OpenStudy (welshfella):

please dont give direct answers EXplain how you got the answer

OpenStudy (bigdaddyharsha):

4^2 + 6^2 = √52 = 7.2ft